# Complements of discriminants of real parabolic function singularities. II

### v4

@@ Abstract.
We list all local connected components of sets of non-discriminant functions near all parabolic function singularities (which are the second most common family of singularity classes of smooth functions after simple singularities). Thus, we prove (and improve in one particular case) all the corresponding conjectures from the previous work [25] with the same title. As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs (one of which is a new one). We also show that the complements of the discriminant varieties of the $X_{9}^{\pm}$ and $P_{8}^{1}$ singularities within their versal deformations have nontrivial one-dimensional homology groups, in contrast to all simple singularities.
These results are applications of a general method for investigating non-singular perturbations of real function singularities. An important part of this method is a computer program that formalizes Picard–Lefschetz theory and surgeries of Morse functions.
@@ 1. Introduction
$1.$
Discriminant varieties of families of smooth functions depending on parameters are the sets of parameters for which the zero-level sets of the corresponding functions are singular. These varieties appear in many problems of PDE theory and integral geometry, often under the name “wavefronts”, see e.g. [2], [3], [4], [9], [14], [21], [23], [24].
Many functions studied in these theories are defined by integral transforms. Their arguments are the parameters of certain function families (called generating families), and their values are defined by the integrals along the cycles in the zero-level sets of the corresponding functions. These integral functions are regular outside the discriminant varieties of the families. The analytic behavior of such a function near a discriminant point depends on the singularity type of the corresponding function and on the connected component of the complement of the discriminant, from which we approach this point. Thus, enumerating these components is one of the first steps in studying the analytical properties of the integral transforms.
@@
$f(x_{1},\dots,x_{n})+Q_{2}(y_{1},\dots,y_{N-k})\sim\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})+\tilde{Q}_{2}(\tilde{y}_{1},\dots,\tilde{y}_{N-m}).$
The critical points of analytic functions ${\mathbb{R}}^{n}\to{\mathbb{R}}$ are equivalent if the function singularities obtained from their germs at these points by subtracting the critical values are equivalent. An $l$-parametric deformation of a function singularity $f$ is a family of analytic functions, $f_{\lambda}$, where $\lambda$ runs through a neighborhood of the origin point $0\in{\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ in $n+l$ variables is regular analytic. The discriminant variety of the deformation $F$ is the set of parameter values $\lambda\in{\mathbb{R}}^{l}$ such that the corresponding function $f_{\lambda}$ has a real critical point with zero critical value near the origin in ${\mathbb{R}}^{n}$ (or, equivalently, the zero-level set of the analytic continuation of $f_{\lambda}$ to ${\mathbb{C}}^{n}$ is singular at this point).
For each isolated function singularity, studying all its deformations reduces to studying any its sufficiently large (so-called versal, see [3]) deformation. The complements of discriminant varieties of all versal deformations of stably equivalent singularities are homotopically equivalent, in particular the connected components of these complements are in a one-to-one correspondence. Therefore, for any stable equivalence class of singularities it is sufficient to enumerate these components for only one versal deformation of one function of this class.
@@
O. Viro [31] found all topological types of non-discriminant perturbations of the $X_{9}$ and $J_{10}$ singularities. By definition, two small non-discriminant perturbations $f_{\lambda}:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of the same isolated function singularity have the same topological type if there is an isotopy between their zero-level sets in a neighborhood of the critical point that preserves the behavior of these sets at the boundary of the neighborhood. Clearly, the topological type is an invariant of the connected components of the discriminant complements. Our “rigid isotopy” classification of non-discriminant perturbations, i.e., the classification whose classes are these components, is only slightly stricter than the topological classification, see Table 2. Counting the components of the discriminant complements of $P_{8}$ singularities is related to (though is different from) the isotopy classification of nonsingular cubic surfaces in ${\mathbb{R}}^{3}$, which is described in detail in [8]. This classification consists of five and ten classes for $P_{8}^{1}$ and $P_{8}^{2}$ singularities, respectively.
@@ 1.2. On the method
In [25], an invariant of the connected components of the complements of the discriminant varieties of function singularities is described, complete lists of possible values of this invariant for parabolic singularities are found, and all of these values are realized by concrete polynomials. The remaining problem is to list all components with each value. In [25], a conjectural answer to this problem was proposed. It provides the lists consisting of 7, 15, 7, 14, 52, 11, and 33 connected components for the singularities of classes $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively. Below, we prove that almost all of these lists are indeed complete. However, some two sets of perturbations of the $J_{10}^{1}$ singularities indicated in [25] consist of pairs of different connected components. Thus, the number 11 is replaced in Table 2 by 13. These results are summarized in Theorems 6, 14, 19, 29, 32, 38, and 42 below.
@@ 1.3. On the discriminant complements for simple singularities
E. Looijenga [16] implicitly listed the local connected components of the complements of the discriminants of simple singularities, although the numbers of these components remained unknown. O. Viro [31] listed all possible topological types of non-discriminant functions representing these components. In [26], it was shown that each topological type on this list corresponds to exactly one connected component of the discriminant complement. For simple singularities with Milnor numbers at most six this fact also follows from the results of [19]. Unfortunately, I only read the work [31], which contains the enumeration of topological types of simple and parabolic singularity perturbations in two variables after the publication of [25] and [26].
@@
E. Looijenga [16] proved that all local components of discriminant complements of real simple singularities are homotopically trivial. Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 13 and 40.
@@
In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [2], [9], [14], [18], [23]). Near the non-singular points of wavefronts (corresponding to the Morse singularities of generating functions) the lacunas were enumerated in [7] and [6], see also [2]. Near the simplest non-Morse points of classes $A_{2}$ and $A_{3}$, this was done in [9], and near all simple singularities in [21]. In [24], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that the list is complete. We prove this conjecture for almost all parabolic classes but discover one additional local lacuna in the $P_{8}^{2}$ case, see § 5.
@@ 2. Results for the $X_{9}^{*}$ classes
Real parabolic function singularities of type $X_{9}$ form four stable equivalence classes $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, and $X_{9}^{2}$.
@@
The parameters $\{\lambda_{1},\dots,\lambda_{8},A\}$ identify the spaces $\Phi_{+}$ and $\Phi_{1}$ with ${\mathbb{R}}^{8}\times(-1,1)$, and the space $\Phi_{2}$ with ${\mathbb{R}}^{8}\times(-\infty,-1)$. The discriminant subvarieties of these spaces consist of polynomials that have at least one real critical point with zero critical value. These varieties divide the spaces $\Phi_{*}$ into connected components. Enumerating these components is our goal in this section.
@@ 2.2. Virtual functions and virtual components
If two generic polynomials, $f$ and $\tilde{f}$, belong to the same connected component of the complement of the discriminant, then the virtual components $S(f)$, S(f~)S($\tilde${f}) are the same. $\Box$
@@ 2.3. Enumeration of components for $X_{9}^{+}$ singularities
Pictures $(f)$ and $(g)$ of Fig. 1 are topologically equivalent. However, Bezout’s theorem arguments show that the polynomials that realize them cannot be connected within the space of non-discriminant polynomials of class $\Phi_{+}$. Figure 1 coincides with the well-known list of affine rigid isotopy classes of bounded degree four smooth algebraic plane curves. In our case, however, the isotopies are allowed only within the space $\Phi_{+}$ which is much smaller than the space of all degree four polynomials. In other cases, $\Phi_{1}$ and $\Phi_{2}$, the analogous restriction leads to an increase of classification lists.
@@
There are exactly seven virtual components of class $\Phi_{+}$. The shapes of the zero-level sets of some polynomials associated with virtual functions representing these virtual components are shown in Fig. 1. $\Box$
@@
Clearly, this group is isomorphic to the group of symmetries of the square, in particular is of order eight. This group acts on the space $\Phi_{+}$ (each transformation $\Gamma:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ moves the polynomial $f$ to $f\circ\Gamma$). This action preserves the set of generic polynomials.
@@
Using the Lyashko–Looijenga covering (in the form described in [12]), we can assume that the functions $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the common associated virtual function. Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has in its interior points only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the union of the discriminant and the set of not strictly Morse functions). Since the virtual function of $\tilde{f}$ is the same as that of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$. (For this existence statement, the parabolicity of $X_{9}$ singularities is essential. The proof of it is based on the results of [12].) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2). First, suppose that it is the same point $I(1)$. A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that preserves the associated virtual functions and commutes with the Lyashko–Looijenga map. This automorphism is induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$, see [28]. Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are either the same or mapped to each other by this automorphism. In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this automorphism, which belongs to the group $G_{0}$. If the point $\tilde{I}(1)$ is different from $I(1)$, we first apply the element of the group $G_{0}$, moving these points to each other, to the entire path $\tilde{I}$ (in particular, to its starting point $\tilde{f}$). This leads us to the situation considered previously. $\Box$
@@
Pictures $(a)$, $(c)$, $(b)$, $(f)$ of Fig. 1 are realized, respectively, by polynomials

### v5

@@ Abstract.
We list all local connected components of sets of non-discriminant functions near all parabolic function singularities (which are the second most important family of singularity classes of smooth functions after simple singularities). Thus, we prove (and improve in one particular case) all the corresponding conjectures from the previous work [25] with the same title. As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs, in particular, find a new one. We also show that the complements of the discriminant varieties of the $X_{9}^{\pm}$ and $P_{8}^{1}$ singularities within their versal deformations have nontrivial one-dimensional homology groups, in contrast to all simple singularities.
These results are applications of a general method for investigating and separating non-singular perturbations of real function singularities. An important part of this method is a computer program that formalizes Picard–Lefschetz theory and surgeries of Morse functions.
@@
This work was supported by the Absorption Center in Science of the Ministry of Immigration and Absorption of the State of Israel
@@ 1. Introduction
1.1. Motivation and terminology
The discriminant variety of a family of smooth functions that depend on parameters is the set of parameters for which the zero-level sets of the corresponding functions are singular. These varieties appear in many problems of PDE theory and integral geometry, often under the name “wavefronts”, see e.g. [2], [3], [4], [9], [14], [21], [23], [24].
Many functions studied in these theories are defined by integral transforms. Their arguments are the parameters of certain function families (called generating families), and their values are defined by the integrals of certain differential forms along the cycles in the zero-level sets of the corresponding functions. These integral functions are regular outside the discriminant varieties of the families. The analytic behavior of such a function near a discriminant point depends on the singularity type of the corresponding function and on the connected component of the complement of the discriminant variety, from which we approach this point. Thus, enumerating these components is one of the first steps in studying the analytical properties of the integral transforms.
@@ 1.1. Motivation and terminology
$f(x_{1},\dots,x_{n})+Q_{2}(y_{1},\dots,y_{N-n})\sim\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})+\tilde{Q}_{2}(\tilde{y}_{1},\dots,\tilde{y}_{N-m})$
for some $N$ and quadratic polynomials $Q_{2},$ $\tilde{Q}_{2}$ of maximal ranks. The critical points of analytic functions ${\mathbb{R}}^{n}\to{\mathbb{R}}$ are equivalent if the function singularities obtained from their germs at these points by subtracting the critical values are equivalent. An $l$-parametric deformation of a function singularity $f$ is a family $\{f_{\lambda}\}$ of analytic functions, where $\lambda$ runs through a neighborhood of the origin point $0\in{\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ is regular analytic in a neighborhood of the origin in ${\mathbb{R}}^{n+l}$. The discriminant variety of the deformation $F$ is the set of parameter values $\lambda\in{\mathbb{R}}^{l}$ such that the corresponding functions $f_{\lambda}$ have real critical points with zero critical value near the origin in ${\mathbb{R}}^{n}$ (or, equivalently, the zero-level sets of the analytic continuations of $f_{\lambda}$ to ${\mathbb{C}}^{n}$ are singular at these points).
For each isolated function singularity, studying all its deformations reduces to studying any its sufficiently large (so-called versal, see [3]) deformation. The complements of discriminant varieties of all versal deformations of stably equivalent singularities are homotopically equivalent, in particular there is a one-to-one correspondence between the connected components of these complements. Therefore, for any stable equivalence class of singularities it is sufficient to list these components for only one versal deformation of one function of this class.
@@
O. Viro [31] found all topological types of non-discriminant perturbations of the $X_{9}$ and $J_{10}$ singularities. By definition, two small non-discriminant perturbations $f_{\lambda}:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of the same isolated function singularity have the same topological type if there is an isotopy between their zero-level sets in a neighborhood of the critical point that preserves the behavior of these sets at the boundary of the neighborhood. Clearly, the topological type is an invariant of the local connected components of the discriminant complements. Our “rigid isotopy” classification of non-discriminant perturbations, i.e., the classification whose classes are these components, is only slightly stricter than the topological classification, see Table 2. Counting the components of the discriminant complements of $P_{8}$ singularities is related to (though is different from) the isotopy classification of nonsingular cubic surfaces in ${\mathbb{R}}^{3}$, which is described in detail in [8]. This classification consists of five and ten classes for $P_{8}^{1}$ and $P_{8}^{2}$ singularities, respectively.
1.2. Structure and interpretation of results
Each class of parabolic singularities is unimodal, that is, it is swept out by a one-parametric family of orbits of the natural action of the group of local diffeomorphisms $({\mathbb{R}}^{n},0)\to({\mathbb{R}}^{n},0)$. In each case, we consider the canonical versal deformation of the corresponding class. This is a space of polynomials that is it diffeomorphic to ${\mathbb{R}}^{9}$ (respectively, ${\mathbb{R}}^{10}$, ${\mathbb{R}}^{8}$) in the case of singularities of classes $X_{9}^{*}$, ($J_{10}^{*}$, $P_{8}^{*}$), see formulas (1), (31), (32), (42). The polynomials with parabolic singularities of the corresponding class form a subset that is diffeomorphic to an open interval in this space. All orbits are represented by some points of these sets. Our main results are the lists of connected components, into which the discriminant varieties divide these spaces, see Theorems 6, 14, 19, 29, 32, 39, and 43. However, each of these results can be interpreted as a list of all local components of the discriminant complements in arbitrarily small neighborhoods of an arbitrary particular parabolic singularity of the corresponding class. Indeed, one-parametric groups of homogeneous or weighted homogeneous dilations act on the parameter spaces of versal deformations. These groups are naturally isomorphic to the multiplicative group ${\mathbb{R}}^{1}_{+}$ of positive numbers. For $P_{8}$ singularities the element $\{t\}\in{\mathbb{R}}^{1}_{+}$ maps each of degree three polynomial $f_{\lambda}(x,y,z)$ to $t^{-3}f_{\lambda}(tx,ty,tz)$, in particular preserves the principal homogeneous part of the polynomials. For $X_{9}$ singularities it acts as $\{t\}:f_{\lambda}(x,y)\to t^{-4}f_{\lambda}(tx,ty)$, and for $J_{10}$ as $\{t\}:f_{\lambda}(x,y)\to t^{-6}f_{\lambda}(t^{2}x,ty)$. All of these actions preserve the discriminants and their complements. Parabolic singularities are the only invariant points of these actions. Each point of the discriminant complement is reduced by this action into an arbitrarily small neighborhood of some parabolic singularity. Additionally, according to [17], the discriminant variety is equisingular along the one-dimensional sets of parabolic singularities in parameter spaces of versal deformations. Therefore, each component of the discriminant complement, which is represented in an arbitrarily small neighborhood of some particular parabolic singularity, is also represented in the neighborhoods of all other parabolic singularities in the one-parametric sets.
@@ 1.3. On the method
In [25], an invariant of the connected components of the complements of the discriminant varieties of function singularities is described, complete lists of possible values of this invariant for parabolic singularities are found, and all of these values are realized by concrete polynomials. The remaining problem is to list all components with each value. In [25], a conjectural answer to this problem was proposed. It provides the lists consisting of 7, 15, 7, 14, 52, 11, and 33 connected components for the singularities of classes $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively. Below, we prove that almost all of these lists are indeed complete. However, some two sets of perturbations of the $J_{10}^{1}$ singularities indicated in [25] consist of pairs of different connected components. Thus, the number 11 is replaced in Table 2 by 13.
@@ 1.4. On the discriminant complements for simple singularities
E. Looijenga [16] implicitly listed the local connected components of the complements of the discriminants of simple singularities, although the numbers of these components remained unknown. O. Viro [31] listed all possible topological types of non-discriminant functions representing these components. In [26], it was shown that each topological type on this list corresponds to exactly one connected component of the discriminant complement. For simple singularities with Milnor numbers at most six this fact also follows from the results of [19]. Unfortunately, I only read the work [31], which contains an enumeration of the topological types of perturbations of simple and parabolic singularities in two variables, after the publication of [25] and [26].
@@
E. Looijenga [16] proved that all local connected components of discriminant complements of real simple singularities are homotopically trivial. Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 13 and 41.
@@
In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [2], [9], [14], [23]). Near the non-singular points of wavefronts (corresponding to the Morse singularities of generating functions) the local lacunas were enumerated in [7] and [6], see also [2]. Near the simplest non-Morse points of classes $A_{2}$ and $A_{3}$, this was done in [9], and near all simple singularities in [21]. In [24], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that this list is complete. We prove this conjecture for almost all parabolic classes but discover one additional local lacuna in the $P_{8}^{2}$ case, see § 5.
A notation
The symbol $\Box$ indicates either the conclusion of a proof or the absence of a proof, as in the case of immediate corollaries or statements supplied with references to the works where they are proven.
@@ 2. Results for the $X_{9}^{*}$ classes
Real parabolic function singularities of type $X_{9}$ form four one-parametric families of stable equivalence classes $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, and $X_{9}^{2}$, see Table 1.
@@
The parameters $\{\lambda_{1},\dots,\lambda_{8},A\}$ identify the spaces $\Phi_{+}$ and $\Phi_{1}$ with ${\mathbb{R}}^{8}\times(-1,1)$, and the space $\Phi_{2}$ with ${\mathbb{R}}^{8}\times(-\infty,-1)$. The discriminant subvarieties divide the spaces $\Phi_{*}$ into connected components. Enumerating these components is our goal in this section.
@@ 2.2. Virtual functions and virtual components
If two generic polynomials, $f$ and $\tilde{f}$, belong to the same connected component of the complement of the discriminant, then the virtual components $S(f)$ and S(f~)S($\tilde${f}) are the same. $\Box$
@@ 2.3. Enumeration of components for $X_{9}^{+}$ singularities
Pictures $(f)$ and $(g)$ of Fig. 1 are topologically equivalent. However, Bezout’s theorem arguments show that the polynomials that realize them cannot be connected within the space of non-discriminant polynomials of class $\Phi_{+}$. Figure 1 essentially coincides with the well-known list of affine rigid isotopy classes of bounded degree four smooth algebraic plane curves. In our case, however, the isotopies are allowed only within the space $\Phi_{+}$ which is much smaller than the space of all degree four polynomials. In other cases, $\Phi_{1}$ and $\Phi_{2}$, the analogous restriction leads to an increase of classification lists.
@@
There are exactly seven virtual components of class $\Phi_{+}$. The shapes of the zero-level sets of some polynomials of this class that are associated with virtual functions representing these virtual components are shown in Fig. 1. $\Box$
@@
Clearly, this group is isomorphic to the group of symmetries of the square, in particular is of order eight. This group acts on the space $\Phi_{+}$ (each transformation $\Gamma:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ moves the polynomial $f$ to $f\circ\Gamma$). This action preserves the set of generic polynomials and the set-valued invariant.
@@
Using the Lyashko–Looijenga covering (in the form described in [12]), we can assume that the polynomials $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the common associated virtual function. Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has in its interior points only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the union of the discriminant and the set of polynomials with fewer than nine critical values). Since the virtual function of $\tilde{f}$ is the same as that of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$. (For this existence statement, the parabolicity of $X_{9}$ singularities is essential. The proof of it is based on Proposition 2 of [12].) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2). First, suppose that it is the same point $I(1)$. A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that preserves the associated virtual functions and commutes with the Lyashko–Looijenga map. This automorphism is induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$, see [28]. Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are either the same or mapped to each other by this automorphism. In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this automorphism, which belongs to the group $G_{0}$. If the point $\tilde{I}(1)$ is different from $I(1)$, we first apply the element of the group $G_{0}$, moving these points to each other, to the entire path $\tilde{I}$ (in particular, to its starting point $\tilde{f}$). This leads us to the situation considered previously. $\Box$
@@
Pictures $(a)$, $(c)$, $(b)$, and $(f)$ of Fig. 1 are realized, respectively, by polynomials